# LLM-based multi-agent poetry generation in non-cooperative environments

### v1

@@ 1 Introduction
The current process of generating poetry in machine learning and NLP differs substantially from the way in which humans learn and compose poetry. Different from the paradigm of training one single poetry generation model that learns from particular datasets, human beings learn within a social context where they interact with others through communication, be it spoken or written, formal or non-formal Jarvis (2012). This raises the question of whether a more human-like generation approach could improve poetry production. Multi-agent systems, with rich applications in social simulation Park et al. (2023); Chuang et al. (2023), are suitable for our goal where human behavior, such as poetry composition, can be effectively replicated through LLM-based agents that can be situated well in social networks. Furthermore, as one of the most fundamental elements for creative composition, "divergent" thinking ability (to deviate from established norms) during the learning phase is crucial for both human and computational creativity (Elgammal et al., 2017; Brinkmann et al., 2023; Wingström et al., 2023). This serves as a rationale for us to adopt a social learning process that supports not only collaboration but also opposition as the main source of deviation in the process Eger (2016); Shi et al. (2019). Moreover, unlike previous studies that focus on weak forms of opposition, such as debates and arguments, which emphasize refining differences through thought correction Chan et al. (2023), our proposed social learning process aims to amplify divergence to enhance diversity.
@@ 2 Related work
The interaction of LLM-based agents The form of interactions among agents can be broadly categorized as cooperative and non-cooperative. Very often, agents communicate cooperatively where the aim is to make joint decisions through collaboration such as back-and-forth communication Li et al. (2024), majority voting Hamilton (2023) or a combination of both Zhuge et al. (2023). Non-cooperative interactions, though less prevalent in comparison, can enhance the quality of responses through debates or arguments among agents Chan et al. (2023); Du et al. (2023). Moreover, while most studies focus on static interaction among agents, some researchers delve into the dynamics of agents’ interactions. (Liu et al., 2023) propose a dynamic interaction architecture where they utilize an optimization algorithm to select the best agents at inference time. Autogen Wu et al. (2023) also enables dynamic group chat to guide the flow of interaction among agents during ongoing conversations. Others focus on the output dynamics during the interaction, instead. (Chuang et al., 2023) find that LLMs tend to align with factual information regardless of their personas and initial states which limits the simulation of opinion dynamics using LLM-based agents. More recently, the dynamics of group interaction is also studied to show that incorporating chain-of-thought reasoning, detailed personas, and finetuning LLMs can enhance agents’ ability to replicate human-like group dynamics Chuang et al. (2024). Our work differs in that 1) we build a social network with group affiliations to obtain a more human-like learning process; 2) we propose a framework that involves two forms of interaction, especially with an emphasis on non-cooperative communication ; 3) we focus on the output dynamics of the generation process under finetuning for training-based agents and consecutive prompting using detailed personas for prompting-based agents.
@@
LM ensemble can be divided into 1) conversational ensemble which does not involve parameter training Wang et al. (2022a) and 2) finetuning-based ensemble i.e., neural network ensemble Shazeer et al. (2016) and output ensemble Dekoninck et al. (2023); Jiang et al. (2023). Prompting-based conversational ensemble is often utilized for reasoning tasks where LLMs ensemble their own responses (i.e., self-ensemble) Wang et al. (2022a); Fu et al. (2022). Very recently, (Lu et al., 2024) combine multiple small conversational models in a parameter-efficient and interpretable way that outperforms ChatGPT according to their A/B test. Finetuning-based ensembles can operate at the neural network or output level. While neural network ensemble typically requires massive training or finetuning through extensive datasets and resources Shazeer et al. (2016); Jiang et al. (2024), the mixture of smaller modules such as adapters becomes a viable solution in resource-restricted situations Wang et al. (2022b); Chronopoulou et al. (2023).
@@
CTG, the task of generating texts subject to attributes such as emotion Firdaus et al. (2022); Ruan and Ling (2021), topic Dathathri et al. (2019); Wang et al. (2019), toxicity avoidance Liu et al. (2021), style Belouadi and Eger (2023); Shao et al. (2021b), debiasing Dinan et al. (2020); Sheng et al. (2020), etc., is also relevant to our study. CTG can operate at the training/finetuning and inference stage. Similarly to LM ensemble, finetuning with additional modules such as task-related adapters Ribeiro et al. (2021); Lin et al. (2021) is also utilized to gain parameter-efficient controllability. Moreover, for CTG applications to reduce the probability of generating undesirable attributes, in addition to the standard Cross-Entropy loss utilized for text generation finetuning, other loss functions are also explored. (Qian et al., 2022) additionally include Contrastive loss which is crucial for the detoxification task but only partially improves the performance of the sentiment control task. (Zheng et al., 2023) also employ a Contrastive loss on sequence likelihood to decrease the generation probability of negative samples. In comparison, operation at the inference stage is more viable in the era of LLMs Jiang et al. (2023); Wang et al. (2023b); Dekoninck et al. (2023). Reranking the outputs is a popular solution. For example, (Jiang et al., 2023) first rerank the complete candidate outputs from multiple LLMs and then fuse the top-K answers. Reranking the original next-token distribution during the decoding stage is also widely explored such as utilizing discriminators Dathathri et al. (2019) or combining opinions (i.e., output logits) from (anti-)expert models Liu et al. (2021); Dekoninck et al. (2023). Operation during inference, especially at the decoding stage, offers strong controllability over the generated texts with less requirement on time and computational resources. However, it may cause a slight decrease in text quality Dathathri et al. (2019). On the other hand, operations at the training/finetuning stage preserve high-quality text with weaker controllability Zhang et al. (2023).
@@
Automatic poetry generation Early attempts to automatic poetry generation mainly rely on grammatical rules Oliveira (2012), statistical rules Jiang and Zhou (2008); Greene et al. (2010) and neural networks such RNNs Zhang and Lapata (2014); Ghazvininejad et al. (2017); Wöckener et al. (2021), especially RNN-based encoder-decoder architecture Wang et al. (2016); Lau et al. (2018); Yan (2016). More recent models focus on transformer-based architectures Tian et al. (2021); Shao et al. (2021a). Although variants of GPT models have demonstrated outstanding performance in many NLP tasks, mixed opinions are observed on their ability to generate poetry. Studies such as (Bena and Kalita, 2019; Liao et al., 2019; LC, 2022) finetune GPT-2 with additional components such as emotion, form and theme and result in moderate to high-quality poems according to human evaluation. (Köbis and Mossink, 2021) claim that zero-shot GPT-2 can generate human-like poems where the best poem according to human selection can match human-written ones but without human preselection, machine-generated poems are easily identifiable. (Wöckener et al., 2021) point out that similar to RNN-based models, GPT-2 faces problems learning poetry-specific characteristics such as rhyme and meter. To counter such deficiencies, (Belouadi and Eger, 2023) propose ByGPT5, an end-to-end token-free model conditioned on rhyme, meter and alliteration. The model can outperform larger models such as GPT-2, ByT5, and ChatGPT (GPT3-3.5) according to both automatic and human evaluation. They also construct a customized corpus QuaTrain consisting of large-scale machine-labeled pseudo-quatrains to enlarge the finetuning dataset size. Moreover, (Sawicki et al., 2023b) claim that GPT-3 finetuned on 300 poems can successfully generate high-quality poems in a specific author’s style but GPT-3.5 without finetuning leads to undesirable poems written with similar styles. The findings from (Sawicki et al., 2023a) also point out that GPT-3.5 and GPT-4, without finetuning, fail to generate poetry in desired styles. Recently, interactive poetry generation has attracted the attention of researchers as it facilitates human-machine collaboration to generate poetry of more diverse styles and better quality under specific constraints Zhipeng et al. (2019); Uthus et al. (2022). (Ma et al., 2023) propose a post-polishing system that fine-grains GPT-2 generation based on constraints from humans. CoPoet from (Chakrabarty et al., 2022) finetunes pretrained T5 with <instruction, generation> pairs to enable poetry generation according to human instructions. Their study shows that finetuned T5 model is not only competitive to the larger InstructGPT model but also successfully collaborates with humans to produce better poems. In our study, we utilize GPT-2 as our base model as it is a good trade-off between parameter efficiency and language proficiency. Contrary to most poetry generation objectives that optimize few poetry-specific characteristics, we leverage poems with arbitrary styles where we initialize our models with randomly drawn samples from QuaTrain corpus that contain pseudo poetic features. Additionally, we also explore the potential of GPT-3.5 and GPT-4 in an interactive environment enabled by a multi-agent system.
@@ 3.3.1 Training-based agents
  • Joint learning with Contrastive loss: Our social network design well suits the setting for Contrastive learning where poems from ‘in-group’ agents are positive samples and poems from ‘out-group’ agents are negative samples. We adopt Contrastive learning to pull closer the semantic representation of ‘in-group’ samples and push apart that of ‘out-group’ samples. We implement the Contrastive loss SimCSE proposed by (Gao et al., 2021). For a mini-batch containing samples from $\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$, let $(o_{j}^{\mathcal{A}_{i}},o_{j}^{\mathcal{A}^{+}},o_{j}^{\mathcal{A}^{-}})$ denote the $j^{th}$ paired triple and $(\boldsymbol{h_{j}},\boldsymbol{h_{j}}^{+},\boldsymbol{h_{j}}^{-})$ be its representation. The Contrastive loss is thus
@@ 4.3 Evaluation
Metric for lexical diversity and novelty. We measure the lexical diversity using the percentage of distinct uni-grams (distinct-1) and bi-grams (distinct-2) following the definition by (Su et al., 2022; Tevet and Berant, 2021). The formulation is given as: unique $\frac{\text{unique n-grams}(O)}{\text{total n-grams}(O)}$, where $O$ is the set of generated poems to be evaluated. We adopt the measure of novelty (novelty-1 and novelty-2) by (McCoy et al., 2023; Shen et al., 2020) where we calculate the number of new uni-/bi-grams that do not stem from the pretraining set and rescale them with the total number of generated tokens. Thus, novelty reflects the lexical difference between the generated poems and the training set, while diversity indicates the token variety among the generated poems.
@@ 5.1.2 Group divergence in semantics
Undesirable increasing semantic similarity from ‘out-group’ agents Figure 5 shows the group divergence of prompting-based agents measured by the mean pairwise semantic similarity over varying prompting strategies and base models. The solid line represents the semantic similarity measured for ‘in-group’ agents and the dashed line for ‘out-group’ agents. We observe an increasing similarity for both ‘in-group’ and ‘out-group’ agents where the greatest group divergence is observed at $t=1$. Over time, the agents tend to generate semantically similar poetry for both GPT-3 and GPT-4. Moreover, we also notice that prompting-based agents generate poetry of homogeneous styles over time which coincides the finding of (Sawicki et al., 2023a). We discuss this point in more detail in Section 5.2.
@@ 6.3 Can different initializations lead to group-based behaviors for prompting-based agents?
As discussed in Section 5, the framework built with prompting-based agents does not exhibit any group-based behavior as expected. Considering that we initialize the agents with random samples drawn from the QuaTrain corpus, we suspect this may cause a high resemblance among the initialized poems. To examine whether an initialization with poems of more contrastive forms can produce group-based behavior, we conduct an experiment using GPT-3.5 under chain-prompting strategy where we initialize group A with poems written by Edgar Allan Poe and group B with poems written by school children under 12 years old Hipson and Mohammad (2020). An example poem from Edgar Allan Poe is “From the lightning in the sky, As it passed me flying by, From the thunder and the storm, And the cloud that took the form.” and an example poem from a school child is “Roses are red, violets are blue. I love the zoo. do you?” We implement the same process and compute the statistics. Slightly surprisingly, we observe a very similar trend for both diversity and semantic divergence to that of random initialized prompting-based agents as shown in Section 5. In terms of diversity, we notice an increase of 2 pp at iteration $t=1$ and then a decreasing trend for both distinct-1 and distinct-2. Qualitatively, at iteration $t=1$, we obtain poems from group B such as “As the sun rose, a butterfly landed softly on my hand, whispering secrets of the garden with each flutter of its delicate wings.”, which resembles the tone of a child and the imagery of a child playing in the garden. However, as $t>1$, we yield similar homogeneous poems to the case in Table 7. An example poem from group B at $t=4$ is “Beneath the starlit sky, a solitary figure stands, A soft whisper of wind caresses the quiet lands. Burdened with untold sorrows in the night so still, ‘I am but a fleeting shadow, lost in time’s skill.’ ”. The results again suggest that GPT-3.5 (also GPT-4) tends to ignore the prompts (i.e., in our case their personas) and rely more on its pretraining knowledge. This observation coincides with the work from (Chuang et al., 2023; Tirumala et al., 2022) that larger models suffer more from memorization.

### v2

@@ 1 Introduction
The current process of generating poetry in machine learning and NLP differs substantially from the way in which humans learn and compose poetry. Different from the paradigm of training one single poetry generation model that learns from particular datasets, human beings learn within a social context where they interact with others through communication, be it spoken or written, formal or non-formal Jarvis (2012). This raises the question of whether a more human-like generation approach could improve poetry production. Multi-agent systems, with rich applications in social simulation Park et al. (2023); Chuang et al. (2023), are suitable for our goal where human behavior, such as poetry composition, can be effectively replicated through LLM-based agents that can be situated well in social networks. Furthermore, as one of the most fundamental elements for creative composition, "divergent" thinking ability (to deviate from established norms) during the learning phase is crucial for both human and computational creativity Elgammal et al. (2017); Brinkmann et al. (2023); Wingström et al. (2023). This serves as a rationale for us to adopt a social learning process that supports not only collaboration but also opposition as the main source of deviation in the process Eger (2016); Shi et al. (2019). Moreover, unlike previous studies that focus on weak forms of opposition, such as debates and arguments, which emphasize refining differences through thought correction Chan et al. (2023), our proposed social learning process aims to amplify divergence to enhance diversity.
@@ 2 Related work
The interaction of LLM-based agents The form of interactions among agents can be broadly categorized as cooperative and non-cooperative. Very often, agents communicate cooperatively where the aim is to make joint decisions through collaboration such as back-and-forth communication Li et al. (2024), majority voting Hamilton (2023) or a combination of both Zhuge et al. (2023). Non-cooperative interactions, though less prevalent in comparison, can enhance the quality of responses through debates or arguments among agents Chan et al. (2023); Du et al. (2023). Moreover, while most studies focus on static interaction among agents, some researchers delve into the dynamics of agents’ interactions. Liu et al. (2023) propose a dynamic interaction architecture where they utilize an optimization algorithm to select the best agents at inference time. Autogen Wu et al. (2023) also enables dynamic group chat to guide the flow of interaction among agents during ongoing conversations. Others focus on the output dynamics during the interaction, instead. Chuang et al. (2023) find that LLMs tend to align with factual information regardless of their personas and initial states which limits the simulation of opinion dynamics using LLM-based agents. More recently, the dynamics of group interaction is also studied to show that incorporating chain-of-thought reasoning, detailed personas, and finetuning LLMs can enhance agents’ ability to replicate human-like group dynamics Chuang et al. (2024). Our work differs in that 1) we build a social network with group affiliations to obtain a more human-like learning process; 2) we propose a framework that involves two forms of interaction, especially with an emphasis on non-cooperative communication ; 3) we focus on the output dynamics of the generation process under finetuning for training-based agents and consecutive prompting using detailed personas for prompting-based agents.
@@
LM ensemble can be divided into 1) conversational ensemble which does not involve parameter training Wang et al. (2022a) and 2) finetuning-based ensemble i.e., neural network ensemble Shazeer et al. (2016) and output ensemble Dekoninck et al. (2023); Jiang et al. (2023). Prompting-based conversational ensemble is often utilized for reasoning tasks where LLMs ensemble their own responses (i.e., self-ensemble) Wang et al. (2022a); Fu et al. (2022). Very recently, Lu et al. (2024) combine multiple small conversational models in a parameter-efficient and interpretable way that outperforms ChatGPT according to their A/B test. Finetuning-based ensembles can operate at the neural network or output level. While neural network ensemble typically requires massive training or finetuning through extensive datasets and resources Shazeer et al. (2016); Jiang et al. (2024), the mixture of smaller modules such as adapters becomes a viable solution in resource-restricted situations Wang et al. (2022b); Chronopoulou et al. (2023).
@@
CTG, the task of generating texts subject to attributes such as emotion Firdaus et al. (2022); Ruan and Ling (2021), topic Dathathri et al. (2019); Wang et al. (2019), toxicity avoidance Liu et al. (2021), style Belouadi and Eger (2023); Shao et al. (2021b), debiasing Dinan et al. (2020); Sheng et al. (2020), etc., is also relevant to our study. CTG can operate at the training/finetuning and inference stage. Similarly to LM ensemble, finetuning with additional modules such as task-related adapters Ribeiro et al. (2021); Lin et al. (2021) is also utilized to gain parameter-efficient controllability. Moreover, for CTG applications to reduce the probability of generating undesirable attributes, in addition to the standard Cross-Entropy loss utilized for text generation finetuning, other loss functions are also explored. Qian et al. (2022) additionally include Contrastive loss which is crucial for the detoxification task but only partially improves the performance of the sentiment control task. Zheng et al. (2023) also employ a Contrastive loss on sequence likelihood to decrease the generation probability of negative samples. In comparison, operation at the inference stage is more viable in the era of LLMs Jiang et al. (2023); Wang et al. (2023b); Dekoninck et al. (2023). Reranking the outputs is a popular solution. For example, Jiang et al. (2023) first rerank the complete candidate outputs from multiple LLMs and then fuse the top-K answers. Reranking the original next-token distribution during the decoding stage is also widely explored such as utilizing discriminators Dathathri et al. (2019) or combining opinions (i.e., output logits) from (anti-)expert models Liu et al. (2021); Dekoninck et al. (2023). Operation during inference, especially at the decoding stage, offers strong controllability over the generated texts with less requirement on time and computational resources. However, it may cause a slight decrease in text quality Dathathri et al. (2019). On the other hand, operations at the training/finetuning stage preserve high-quality text with weaker controllability Zhang et al. (2023).
@@
Automatic poetry generation Early attempts to automatic poetry generation mainly rely on grammatical rules Oliveira (2012), statistical rules Jiang and Zhou (2008); Greene et al. (2010) and neural networks such RNNs Zhang and Lapata (2014); Ghazvininejad et al. (2017); Wöckener et al. (2021), especially RNN-based encoder-decoder architecture Wang et al. (2016); Lau et al. (2018); Yan (2016). More recent models focus on transformer-based architectures Tian et al. (2021); Shao et al. (2021a). Although variants of GPT models have demonstrated outstanding performance in many NLP tasks, mixed opinions are observed on their ability to generate poetry. Studies such as Bena and Kalita (2019); Liao et al. (2019); LC (2022) finetune GPT-2 with additional components such as emotion, form and theme and result in moderate to high-quality poems according to human evaluation. Köbis and Mossink (2021) claim that zero-shot GPT-2 can generate human-like poems where the best poem according to human selection can match human-written ones but without human preselection, machine-generated poems are easily identifiable. Wöckener et al. (2021) point out that similar to RNN-based models, GPT-2 faces problems learning poetry-specific characteristics such as rhyme and meter. To counter such deficiencies, Belouadi and Eger (2023) propose ByGPT5, an end-to-end token-free model conditioned on rhyme, meter and alliteration. The model can outperform larger models such as GPT-2, ByT5, and ChatGPT (GPT3-3.5) according to both automatic and human evaluation. They also construct a customized corpus QuaTrain consisting of large-scale machine-labeled pseudo-quatrains to enlarge the finetuning dataset size. Moreover, Sawicki et al. (2023b) claim that GPT-3 finetuned on 300 poems can successfully generate high-quality poems in a specific author’s style but GPT-3.5 without finetuning leads to undesirable poems written with similar styles. The findings from Sawicki et al. (2023a) also point out that GPT-3.5 and GPT-4, without finetuning, fail to generate poetry in desired styles. Recently, interactive poetry generation has attracted the attention of researchers as it facilitates human-machine collaboration to generate poetry of more diverse styles and better quality under specific constraints Zhipeng et al. (2019); Uthus et al. (2022). Ma et al. (2023) propose a post-polishing system that fine-grains GPT-2 generation based on constraints from humans. CoPoet from Chakrabarty et al. (2022) finetunes pretrained T5 with <instruction, generation> pairs to enable poetry generation according to human instructions. Their study shows that finetuned T5 model is not only competitive to the larger InstructGPT model but also successfully collaborates with humans to produce better poems. In our study, we utilize GPT-2 as our base model as it is a good trade-off between parameter efficiency and language proficiency. Contrary to most poetry generation objectives that optimize few poetry-specific characteristics, we leverage poems with arbitrary styles where we initialize our models with randomly drawn samples from QuaTrain corpus that contain pseudo poetic features. Additionally, we also explore the potential of GPT-3.5 and GPT-4 in an interactive environment enabled by a multi-agent system.
@@ 3.3.1 Training-based agents
  • Joint learning with Contrastive loss: Our social network design well suits the setting for Contrastive learning where poems from ‘in-group’ agents are positive samples and poems from ‘out-group’ agents are negative samples. We adopt Contrastive learning to pull closer the semantic representation of ‘in-group’ samples and push apart that of ‘out-group’ samples. We implement the Contrastive loss SimCSE proposed by Gao et al. (2021). For a mini-batch containing samples from $\mathcal{A}_{i},\mathcal{A}^{+},\mathcal{A}^{-}$, let $(o_{j}^{\mathcal{A}_{i}},o_{j}^{\mathcal{A}^{+}},o_{j}^{\mathcal{A}^{-}})$ denote the $j^{th}$ paired triple and $(\boldsymbol{h_{j}},\boldsymbol{h_{j}}^{+},\boldsymbol{h_{j}}^{-})$ be its representation. The Contrastive loss is thus
@@ 4.3 Evaluation
Metric for lexical diversity and novelty. We measure the lexical diversity using the percentage of distinct uni-grams (distinct-1) and bi-grams (distinct-2) following the definition by Su et al. (2022); Tevet and Berant (2021). The formulation is given as: unique $\frac{\text{unique n-grams}(O)}{\text{total n-grams}(O)}$, where $O$ is the set of generated poems to be evaluated. We adopt the measure of novelty (novelty-1 and novelty-2) by McCoy et al. (2023); Shen et al. (2020) where we calculate the number of new uni-/bi-grams that do not stem from the pretraining set and rescale them with the total number of generated tokens. Thus, novelty reflects the lexical difference between the generated poems and the training set, while diversity indicates the token variety among the generated poems.
@@ 5.1.2 Group divergence in semantics
Undesirable increasing semantic similarity from ‘out-group’ agents Figure 5 shows the group divergence of prompting-based agents measured by the mean pairwise semantic similarity over varying prompting strategies and base models. The solid line represents the semantic similarity measured for ‘in-group’ agents and the dashed line for ‘out-group’ agents. We observe an increasing similarity for both ‘in-group’ and ‘out-group’ agents where the greatest group divergence is observed at $t=1$. Over time, the agents tend to generate semantically similar poetry for both GPT-3 and GPT-4. Moreover, we also notice that prompting-based agents generate poetry of homogeneous styles over time which coincides the finding of Sawicki et al. (2023a). We discuss this point in more detail in Section 5.2.
@@ 6.3 Can different initializations lead to group-based behaviors for prompting-based agents?
As discussed in Section 5, the framework built with prompting-based agents does not exhibit any group-based behavior as expected. Considering that we initialize the agents with random samples drawn from the QuaTrain corpus, we suspect this may cause a high resemblance among the initialized poems. To examine whether an initialization with poems of more contrastive forms can produce group-based behavior, we conduct an experiment using GPT-3.5 under chain-prompting strategy where we initialize group A with poems written by Edgar Allan Poe and group B with poems written by school children under 12 years old Hipson and Mohammad (2020). An example poem from Edgar Allan Poe is “From the lightning in the sky, As it passed me flying by, From the thunder and the storm, And the cloud that took the form.” and an example poem from a school child is “Roses are red, violets are blue. I love the zoo. do you?” We implement the same process and compute the statistics. Slightly surprisingly, we observe a very similar trend for both diversity and semantic divergence to that of random initialized prompting-based agents as shown in Section 5. In terms of diversity, we notice an increase of 2 pp at iteration $t=1$ and then a decreasing trend for both distinct-1 and distinct-2. Qualitatively, at iteration $t=1$, we obtain poems from group B such as “As the sun rose, a butterfly landed softly on my hand, whispering secrets of the garden with each flutter of its delicate wings.”, which resembles the tone of a child and the imagery of a child playing in the garden. However, as $t>1$, we yield similar homogeneous poems to the case in Table 7. An example poem from group B at $t=4$ is “Beneath the starlit sky, a solitary figure stands, A soft whisper of wind caresses the quiet lands. Burdened with untold sorrows in the night so still, ‘I am but a fleeting shadow, lost in time’s skill.’ ”. The results again suggest that GPT-3.5 (also GPT-4) tends to ignore the prompts (i.e., in our case their personas) and rely more on its pretraining knowledge. This observation coincides with the work from Chuang et al. (2023); Tirumala et al. (2022) that larger models suffer more from memorization.